# Leveraging Robust Optimization for LLM Alignment under Distribution Shifts

### v3

@@ 1 Introduction
The rapid proliferation of large language models (LLMs) has made it increasingly important to ensure that model outputs align with human values. Techniques such as Reinforcement Learning from Human Feedback (RLHF) [32, 50, 39] and Direct Preference Optimization (DPO) [34] have shown promise by leveraging high-quality, human-annotated data to guide model behavior [42, 2, 13]. A common approach to constructing alignment data involves either manual annotation by humans or the use of LLMs, with the former being resource-intensive and time-consuming, thus limiting scalability and broader applicability [4]. Consequently, recent research has increasingly focused on leveraging synthetic data generated by advanced LLMs, which have shown strong capability to simulate human preferences [23, 7, 8].
@@
[…]
@@ 2 Problem Formulation
Notations. Let $\mathbf{z}=\{x,y_{0},\dots,y_{n-1}\}\in Z$ represent a datum in a preference dataset $Z$, where each instance composes of one query $x$ and $n$ corresponding responses $y_{0},y_{1},y_{2},\dots,y_{n-1}$ We assume that these observed data are drawn from a training distribution $P$, and the target distribution of interest, is denoted as $Q_{0}$. We then formalize the following definition.
@@
Distributionally robust optimization: Distributionally Robust Optimization (DRO) [3, 12] is a well-established framework that focuses on minimizing the worst-case expected loss over the perturbed or adversarial distribution $Q$ within the uncertainty set of $P$. Denote the instance-level loss as $\ell(\theta,\mathbf{z})$, then DRO is formulated as:
@@ 3.1 Sample-Level Calibration with Probabilistic Classifiers
where $i\in\{1,\dots,n-1\}$ and $\gamma_{i}$ is the ratio of the sizes of the two sub-distributions. In practice, to estimate $P_{\text{golden}}$, we leverage human-preferred responses from the target distribution $Q_{0}$.
@@ 3.2 Deriving the DoRA objective
[…]
@@
4:     Sample batch =$\mathcal{B}=\{\mathbf{z}_{i}\}_{i=1}^{b}\subset\mathcal{D}$
@@ 4.1 Experimental Results
Table 1: Comparison with pairwise robust baselines on HH dataset. While all the approaches improve the win rate of vanilla DPO, DoRA achieves the lowest lose rates compared to the counterparts.
@@
Alignment HH-RLHF Summarization
@@
Win(%$\uparrow$ Lose(%$\downarrow$ Win(%$\uparrow$ Lose(%$\downarrow$ Win(%$\uparrow$ Lose(%$\downarrow$ Win(%$\uparrow$ Lose(%$\downarrow$
@@
Methods AlpacaEval 2.0 Arena-Hard
LC(%) WR(%) Length WR(%)
@@ 4.2 Ablations
Figure 6: Impact of different $\lambda$ values on policy distribution. (a). Larger $\lambda$ increases model probabilities on the HH-chosen responses. (b/c). Scatter plots comparing log probabilities assigned to HH-preferred responses. The x-axis shows log probabilities under the initial SFT policy; the y-axis shows those under the original baseline and DoRA policies, respectively. As $\lambda$ increases, DoRA shifts closer to the original baseline (ERM training), assigning higher likelihoods to preferred responses.
@@ 5 Related Works
Preference alignment typically relies heavily on human-annotated datasets, but the high cost and limited scalability of such data present a major bottleneck. To address this, recent research has explored leveraging synthetic data for alignment. For example, RLAIF [23] synthesizes preference data and uses PaLM 2 for feedback, while UltraFeedback [7] employs GPT-4 to annotate LLM-generated responses, creating scalable training datasets. Moreover, researchers have integrated synthetic data to expand candidate pools for preference learning [38, 47, 29, 49], demonstrating the potential of synthetic data. However, the use of synthetic data introduces new challenges. One key issue is the mismatch between the sampling distribution and the learning policy. To address this, RSO [28] employs rejection sampling to source preference data from the estimated target optimal policy, thereby improving the accuracy of the maximum likelihood estimator. Another critical challenge lies in the distributional inconsistencies between synthetic and human-generated data during preference learning. This shift can hinder alignment performance, leading to biased model behaviors to true human preferences. In this paper, we focus on tackling this latter challenge, aiming to enhance the robustness of preference learning in the presence of distribution shifts.
@@ Conclusion.
Our framework assumes access to coarse-grained information about the source of the data (e.g., human- or model-generated) to guide classifier training. While this is often feasible in curated alignment pipelines, it may be less accessible in fully open-domain or legacy datasets. Exploring ways to relax this requirement through unsupervised or weakly supervised signals presents an exciting direction for future work. Our use of DRO is particularly beneficial in scenarios with distribution shifts or variable alignment quality. While its conservativeness may offer limited gains in clean settings, it holds promise in reliability-critical domains such as factual generation or value-sensitive applications. Finally, DoRA currently operates offline. Future work includes extending it to online settings such as self-play, enabling continual adaptation and scalability.
@@ C.1 Data Generation
Eurus-RM-7b (top 10 for Chat) HH-RLHF 65.0
@@ C.2 Reward models
HH-RLHF $\beta$=0.1 $\beta$=0.1, $\alpha$=0.01 $\beta$=0.5 $\beta$=2.0,$\gamma$=0.8 $\beta$=0.1 $\alpha$=1.0 $T$=2.0
Summarization - - - - $\beta$=0.5 $\alpha$=0.5 $T$=1.0
UltraFeedback - - - - $\beta$=0.1 $\alpha$=1.0 $T$=2.0
@@ C.4 Evaluation prompts
For HH and Summarization tasks, we adapt the evaluation prompts from Rafailov et al. [34] using GPT-4o, and compute the win rates and lose rates with 400 randomly selected test queries, with the order randomly swapped. {mdframed}[linewidth=1pt, linecolor=black, backgroundcolor=gray!10] [HH-RLHF]: For the following query to a chatbot, which response is more helpful?
@@
{mdframed}
[linewidth=1pt, linecolor=black, backgroundcolor=gray!10] [Summarization]: Which of the following summaries does a better job of summarizing the most important points in the given forum post, without including unimportant or irrelevant details? A good summary is both precise and concise.
@@ Appendix E Further discussion on RSO
Baselines Methods HH-RLHF
@@
Table 8: Comparison of RSO and DoRA on HH-RLHF. Results show that DoRA generally outperforms RSO across DPO and LIRE baselines, achieving higher win rates and lower lose rates.

### v4

@@ 1 Introduction
The rapid proliferation of large language models (LLMs) has made it increasingly important to ensure that model outputs align with human values. Techniques such as Reinforcement Learning from Human Feedback (RLHF) [33, 51, 40] and Direct Preference Optimization (DPO) [35, 45] have shown promise by leveraging high-quality, human-annotated data to guide model behavior [43, 2, 12]. A common approach to constructing alignment data involves either manual annotation by humans or the use of LLMs, with the former being resource-intensive and time-consuming, thus limiting scalability and broader applicability [4]. Consequently, recent research has increasingly focused on leveraging synthetic data generated by advanced LLMs, which have shown strong capability to simulate human preferences [22, 7, 8].
@@
‘ Contributions We introduce DoRA, a distribution-aware optimization framework that robustly aligns LLM outputs with human preferences under distribution shifts. Technically: We estimate the alignment of each sample with the target human-preferred distribution as a calibration term, which is then incorporated into a distribution-aware optimization objective that minimizes the worst-case loss over regions most relevant to human preferences. Theoretically: We characterize the robust calibration alignment objective as a KL-divergence-based distributionally robust optimization problem, augmented with a reweighting mechanism. Emperically: We demonstrate the effectiveness of our framework through extensive experiments, with consistent improvements in alignment metrics compared to state-of-the-art baselines.
@@ 2 Problem Formulation
Notations. Let $\mathbf{z}=\{x,y_{0},\dots,y_{n-1}\}\in Z$ represent a datum in a preference dataset $Z$, where each instance composes of one query $x$ and $n$ corresponding responses $y_{0},y_{1},y_{2},\dots,y_{n-1}$ . We assume that these observed data are drawn from a training distribution $P$, and the target distribution of interest, is denoted as $Q_{0}$. We then formalize the following definition.
@@
Distributionally robust optimization: Distributionally Robust Optimization (DRO) [3, 11, 26] is a well-established framework that focuses on minimizing the worst-case expected loss over the perturbed or adversarial distribution $Q$ within the uncertainty set of $P$. Denote the instance-level loss as $\ell(\theta,\mathbf{z})$, then DRO is formulated as:
@@ 3.1 Sample-Level Calibration with Probabilistic Classifiers
where $i\in\{1,\dots,n-1\}$ . $\gamma_{i}$ is the imbalance ratio between a synthetic distribution and the golden distribution $\frac{q(y=0)}{q({y=1})}$. In practice, to estimate $P_{\text{golden}}$, we leverage human-preferred responses from the target distribution $Q_{0}$.
@@ 3.2 Deriving the DoRA objective
$\max$ $\mathbb{E}_{P}[h(\mathbf{z})\tilde{h}(\mathbf{z})\ell(\theta,\mathbf{z})],\quad\text{s.t.}\$$\mathbb{E}_{P}[h(\mathbf{z})\log h(\mathbf{z})]\leq\rho,$ (7)
@@
4:  Sample batch $\mathcal{B}=\{\mathbf{z}_{i}\}_{i=1}^{b}\subset\mathcal{D}$
@@ 4.1 Experimental Results
Table 1: Comparison with pairwise robust baselines on $H$ dataset. While all the approaches improve the win rate of vanilla DPO, DoRA achieves the lowest lose rates compared to the counterparts.
@@
Alignment $H$-RLHF Summarization
@@
[…]
@@
Dataset AlpacaEval 2.0 Arena-Hard
Metric LC(%) WR(%) Length WR(%)
@@ 4.2 Ablations
Figure 6: Impact of different $\lambda$ values on policy distribution. (a). Larger $\lambda$ increases model probabilities on the $H$-chosen responses. (b/c). Scatter plots comparing log probabilities assigned to $H$-preferred responses. The x-axis shows log probabilities under the initial SFT policy; the y-axis shows those under the original baseline and DoRA policies, respectively. As $\lambda$ increases, DoRA shifts closer to the original baseline (ERM training), assigning higher likelihoods to preferred responses.
@@
4.3 Generalization and Robustness Evaluation
Table 5: Robustness evaluation under different corruption rates on HH and AlpacaEval 2 datasets. DoRA consistently improves robustness across methods.
Dataset $H$-RLHF AlpacaEval 2
Corruption rate 20% 40% 60 % 40 % 40 %
Metric WR(%) LC(%) WR(%)
[…]
w/ DoRA 74.5 67.0 60.8 12.1 12.1
RRHF 63.5 42.3 26.3 7.7 4.8
w/ DoRA 65.5 44.5 30.8 8.5 5.5
LIRE 67.7 52.5 61.4 26.5 25.1
w/ DoRA 71.5 55.0 65.3 27.7 26.2
Label noise. In the preceding experiments, all models were trained on clean datasets without synthetic perturbations. To further assess the robustness of DoRA, we conduct corrupted-label experiments that serve as stress tests under adversarial distribution shifts. In particular, label noise is introduced by randomly flipping a proportion of labels in the original training data. Experimental results indicate that the incorporation of DoRA consistently enhances model robustness as the corruption rate increases, leading to stable and significant improvements across multiple benchmarks. This further demonstrates DoRA’s effectiveness and reliability in real-world applications where annotation noise or distributional shifts are inevitable.
Table 6: Performance improvement across training iterations. DoRA consistently enhances final performance under self-training scenario.
Method Iter 1 Iter 2 Iter 3
[…]
w/ DoRA 78.5 85.8 88.0
RRHF 45.8 47.8 50.5
w/ DoRA 47.5 49.3 52.5
LIRE 80.3 83.0 84.5
w/ DoRA 82.5 85.0 86.8
Self-training. While the primary focus of this work is on mixture response shifts induced by synthetic data in offline training, we consider online adaptation as a promising direction. To provide preliminary evidence, we conduct a small-scale experiment where we first train LLaMA-3.2-3B on the HH dataset (Iterate 1), and then continue training it on its own generated responses for two additional iterations. This setup naturally induces an evolving mixture shift, as the response distribution becomes increasingly synthetic. The results demonstrate that DoRA remains robust under such conditions, indicating that its instance-level calibration generalizes well to online or self-generated data scenarios where distributional drift arises organically.
@@ 5 Related Works
Preference alignment typically relies heavily on human-annotated datasets, but the high cost and limited scalability of such data present a major bottleneck. To address this, recent research has explored leveraging synthetic data for alignment. For example, RLAIF [22] synthesizes preference data and uses PaLM 2 for feedback, while UltraFeedback [7] employs GPT-4 to annotate LLM-generated responses, creating scalable training datasets. Moreover, researchers have integrated synthetic data to expand candidate pools for preference learning [39, 48, 30, 50], demonstrating the potential of synthetic data. However, the use of synthetic data introduces new challenges. One key issue is the mismatch between the sampling distribution and the learning policy. To address this, RSO [29] employs rejection sampling to source preference data from the estimated target optimal policy, thereby improving the accuracy of the maximum likelihood estimator. Another critical challenge lies in the distributional inconsistencies between synthetic and human-generated data during preference learning. This shift can hinder alignment performance, leading to biased model behaviors to true human preferences[27]. In this paper, we focus on tackling this latter challenge, aiming to enhance the robustness of preference learning in the presence of distribution shifts.
@@ Conclusion.
Our framework assumes access to coarse-grained information about the source of the data (e.g., human- or model-generated) to guide classifier training. While this is often feasible in curated alignment pipelines, it may be less accessible in fully open-domain or legacy datasets. Moreover, DoRA presumes that the target distribution can be at least approximately estimated or measured; in scenarios where such reference distributions are unavailable or unreliable, its applicability may be limited. Identifying principled heuristics or proxy objectives for such cases represents an important direction for future work. Exploring ways to relax this requirement through unsupervised or weakly supervised signals presents an exciting direction for future research. Our use of DRO is particularly beneficial in scenarios with distribution shifts or variable alignment quality. While its conservativeness may offer limited gains in clean settings, it holds promise in reliability-critical domains such as factual generation or value-sensitive applications. Finally, although online adaptation is briefly mentioned in Section 4.3, extending DoRA to broader online learning scenarios remains a promising direction toward continual adaptation and scalability.
Acknowledgments and Disclosure of Funding
This research is supported by Artificial Intelligence-National Science and Technology Major Project 2023ZD0121200 and the National Science Fund for Excellent Young Scholars under Grant 62222212. Besides, we are grateful to Jiashuo Liu and Junkang Wu for providing helpful discussions during the preparation of this paper.
@@ C.1 Data Generation
Eurus-RM-7b (top 10 for Chat) $H$-RLHF 65.0
@@ C.2 Reward models
[…]
Summarization - - - - $\beta$=0.5 $\alpha$=0.5 TT=1.0
UltraFeedback - - - - $\beta$=0.1 $\alpha$=1.0 TT=2.0
@@ C.4 Evaluation prompts
For HH and Summarization tasks, we adapt the evaluation prompts from Rafailov et al. [35] using GPT-4o, and compute the win rates and lose rates with 400 randomly selected test queries, with the order randomly swapped. [$H$-RLHF]: For the following query to a chatbot, which response is more helpful?
@@
[Summarization]: Which of the following summaries does a better job of summarizing the most important points in the given forum post, without including unimportant or irrelevant details? A good summary is both precise and concise.
@@ Appendix E Further discussion on RSO
Baselines Methods $H$-RLHF
@@
Table 10: Comparison of RSO and DoRA on $H$-RLHF. Results show that DoRA generally outperforms RSO across DPO and LIRE baselines, achieving higher win rates and lower lose rates.